OpenStudy (anonymous):

Segment LM is the midsegment of trapezoid ABCD. mc028-2.jpg What is the value of x?

OpenStudy (anonymous):

Mind melting...... 26 214 34 30

OpenStudy (anonymous):

6 more to go...I am almost there....

OpenStudy (jdoe0001):

I'm thinking you're missing mc028-2.jpg , which has the "x" in it :P

OpenStudy (anonymous):

ABCDE AB=x=8 LM=4x+3 DC=180 x+8=4x+3 x+8-3=4x+3-3 x+5=4x 5=3x I'm lost.....

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

can you help me from where I am?

OpenStudy (anonymous):

ABCDE AB=x=8 LM=4x+3 DC=180 x+8=4x+3 x+8-3=4x+3-3 x+5=4x 5=3x

OpenStudy (jdoe0001):

that picture looks like the other one.... but it's just a trapezoid, there's no "x" in it

OpenStudy (anonymous):

This is the pic for the question I am on?

OpenStudy (anonymous):

AB=x=8 LM=4x+3 Got to solve the x

OpenStudy (anonymous):

ABCDE AB=x=8 LM=4x+3 DC=180 x+8=4x+3 x+8-3=4x+3-3 x+5=4x 5=3x

OpenStudy (jdoe0001):

ohh... ahem, I well

OpenStudy (jdoe0001):

AB = x-8?

OpenStudy (anonymous):

That is okay, I'll figure it out

OpenStudy (anonymous):

my choices 26 214 34 30